A vessel of 80ltrs is filled with milk and water. 70% of milk and


30% of water is taken out of the vessel.it is found that the


vessel is vacated by 55%. find the initial quantity of milk and


water???

Grade:

1 Answers

mahendra shah
10 years ago

milk = x litres

water = y litres..

x + y =80...(1)

now,

30x/100 + 70y/100 = 80*45/100

so 3x + 7y = 360...(2)

solve these 2 equations..

y = 30, x = 50!